# DRL-Based Resource Allocation for Motion Blur Resistant Federated Self-Supervised Learning in IoV

### v1

@@ V DRL-BFSSL Algorithm
The optimization problem $\mathcal{C}$ is a NonLinear Programming (NLP) problem. As the number of the vehicles increases, it is hard to address rapidly with the traditional optimization algorithms. Thus, to achieve the optimal solution, we will consider a DRL offloading process. As we know, the SAC algorithm can deal with continuous variables in non-convex function, and keep robustness and adaptability [47]. Thus, we adopt SAC algorithm to solve the optimization problem $\mathcal{C}$. In SAC algorithm, it consists of five Deep Neural Networks (DNNs), i.e., one actor network $\varsigma$, two critic networks $\xi_{1}$ and $\xi_{2}$, and two target critic networks $\varpi_{1}$ and $\varpi_{2}$. The two critic networks work to enhance network performance by minimizing positive bias. The two target networks have the same structure as the critic networks, aiming to improve training speed and stability. By utilizing reward function and Stochastic Gradient Descent (SGD) algorithm, SAC allows us to address the problem in continuous space [61][62].
@@ V-4 SAC-based solution
BS stores the parameters of a global model $\theta_{r-1}^{{\text{global}}}$ which is aggregated at the end of previous round $r-1$ and estimates the positions according to the uploaded velocities $\left\{v_{k,t}^{1},v_{k,t}^{2},\ldots v_{k,t}^{n}\ldots v_{k,t}^{N}\right\}$ and get state $s_{k,t}$ according to Eq. (33). In put $s_{k,t}$ into actor network and output $a_{k,t}$, including transmission powers and CPU frequencies for vehicles. Then, vehicle $n$ downloads the parameters of the global model $\theta_{r-1}^{{\text{global}}}$, transmission power $p_{k,t}^{n}$ , CPU frequency $f_{k,t}^{n}$ from the BS, and sets the parameter of global model to local model $\theta_{r}^{{\text{local},n}}$. If in the first round, the parameters of global model comes from $\theta_{0}^{{\text{global}}}$. If in the episode $k$, slot $1$, the parameters of SAC networks come from the output of episode $k-1$, slot $S^{\text{max}}$.
@@
Next, the maximum number of iterations $\mathcal{N}_{k,t}^{n}$ of local training for vehicle $n$ is calculated according to Eq. (15) and Eq. (16) based on the downloaded CPU frequency $f_{k,t}^{n}$. Then each training vehicle retrains the local model for $\mathcal{N}_{k,t}^{n}$ iterations. Specifically, for each iteration $\psi_{k,t}^{n}\in[1,\mathcal{N}_{k,t}^{n}]$ of local training, each vehicle $n$ encodes the images of the local data, where the encoding process is illustrated in Fig. 3. For each image $x_{r}^{i,n}\in\left\{x_{r}^{1,n},\ x_{r}^{2,n},\ \ldots x_{r}^{i,n},\ldots x_{r}^{\mathcal{Z},n}\right\}$ in vehicle $n$ undergoes two different data augmentation methods $\pi_{1}(\cdot)$ and $\pi_{2}(\cdot)$, e.g., flipping, rotating, scaling, etc. We treat the remaining images, donated as $x_{r}^{j,n},j\in\left[1,\mathcal{Z}\right]\ \text{and}\ j\neq i$, as negative samples. Then, these two augmented images $\pi_{1}(x_{r}^{i,n})$ and $\pi_{2}(x_{r}^{i,n})$ along with $x_{r}^{j,n}$ are input into an $\chi$ function, which composed of an encoder with function $\mathcal{G}_{r}^{\text{local},n}$, and a project head including two dense neutral networks and a ReLU function. The $\chi$ function outputs anchor sample $q_{r}^{i,{n}}$, positive sample $k_{{r}}^{i,{n}}$, and encoded negative samples $k_{{r}}^{j,{n}}$, respectively, i.e.,
@@
Then, for each image, the aim is to minimize the loss function and get optimize local mode $\hat{\theta}_{{r}}^{{\text{local}},n}$, i.e.,
@@
where $\theta_{{r}}^{{\text{local},n}}$ represents the parameters of local model of vehicle $n$ in round $r$. Each vehicle approaches ${\hat{\theta}}_{{r}}^{{\text{local},n}}$ according to SGD method, and the process of updating is as follows:
@@
According to Eq. (45), we assign smaller weights to the parameters obtained from training vehicles with higher blur levels in the local model, thus reducing the impact of the motion blur and improving the performance of the global model. Then the BS employs a weighted federated algorithm to aggregate the parameters of $N$ models based on the blur level $\mathcal{B}_{r}^{n}$. The expression for the aggregated new global model is
@@
Repeat Step 2 to Step 4 until slot $t$ reaches to $S^{\text{max}}$. At this point, if the episode $k$ is divided by $K_{u}$, the five networks of SAC need to be updated, otherwise, it is not required. After that, episode $k$ moves to episode $k+1$, Slot $t$ reset to 00. Next, we will introduce the update process.
@@
For each iteration of the updating process, randomly sample $\mathcal{M}$ tuples from the replay buffer to form a mini-batch for training. For the $m$-th tuple $(s_{m},a_{m},r_{m},s_{m+1})$ in the mini-batch, where $m\in[1,\mathcal{M}]$, inputting $s_{m}$ to the actor network to obtain $a_{m}^{\prime}$. It is worth noting that this action $a_{m}^{\prime}$ is not the same as $a_{m}$ in the tuple. Thus, the gradient of loss function of $\alpha$ can be calculated as
@@
Function UPDATE($\gamma$, $\alpha$, $\varsigma$, $\xi_{1}$, $\xi_{2}$, $\varpi_{1}$, $\varpi_{2}$, $\mathcal{M}$, $k$, $K_{t}$):
      Randomly sample $\mathcal{M}$ tuples from replay buffer
      Update $\alpha$ according to Eq. (49);
      Update $\varsigma$ according to Eq. (51);
      Update $\xi_{1}$ and $\xi_{2}$ according to Eq. (52) and Eq. (53);
      if $k\mod K_{t}=0$ then
             Update $\varpi_{1}$ and $\varpi_{2}$ according to Eq. (55) and Eq. (56);
@@
Initialize round $r\leftarrow 0$, episode $k\leftarrow 0$;
for episode $k$ from $1$ to $K^{\text{max}}$ do
       slot $t\leftarrow 0$;
       for slot $t$ from $1$ to $S^{\text{max}}$ do
             $r\leftarrow(k-1)\cdot S^{\text{max}}+t$;
             Obtain state $s_{k,t}$ based on distance and velocity;
             Get action $a_{k,t}$ through SAC algorithm;
             for $n=1,2,\dots,N$ in parallel do
                   Download global model $\theta_{r-1}^{\text{global}}$, $p_{k,t}^{n}$, and $f_{k,t}^{n}$;
                   $\theta_{r}^{{\text{local}},n}\leftarrow\text{LOCAL}(\theta_{r-1}^{\text{global}},p_{k,t}^{n},f_{k,t}^{n})$;
            if $r=1$ then
                   Vehicle $n\in[1,N]$ uploads the local model $\theta_{r}^{{\text{local}},n}$ and $v_{k,t}^{n}$ to BS with download transmission power $p_{k,t}^{n}$, along with velocity $v_{0}^{n}$;
            else
                   Vehicle $n\in[1,N]$ uploads the local model $\theta_{r}^{{\text{local}},n}$ and $v_{k,t}^{n}$ to BS with download transmission power $p_{k,t}^{n}$;
            BS aggregates received local models based on $\mathcal{B}_{r}^{n}$ and updates $\theta_{r}^{\text{global}}$ according to Eq. (46);
             Save $\theta_{r}^{\text{global}}$ to global model;
      if episode $k$ mod $K_{u}$ = 0 then
             UPDATE($\gamma$, $\alpha$, $\varsigma$, $\xi_{1}$, $\xi_{2}$, $\varpi_{1}$, $\varpi_{2}$);
Function Local($\theta_{r-1}^{\text{global}},p_{k,t}^{n},f_{k,t}^{n}$):
@@
       Set the parameter of global model $\theta_{r-1}^{\text{global}}$ to local model $\theta_{r}^{{\text{local}},n}$;
       Calculate iteration value $\mathcal{N}_{k,t}^{n}$ according to Eq. (15) and Eq. (16);
       for each iteration $\psi_{k,t}^{n}$ from $1$ to $\mathcal{N}_{k,t}^{n}$ do
             Prepare local model $\theta_{r}^{{\text{local}},n}$;
             for image $i$ from $1$ to $\mathcal{Z}$ do
                   Calculate $q_{r}^{i,n}$, $k_{r}^{i,n}$, and $k_{r}^{j,n}$ according to Eq. (37) - Eq. (39);
            Calculate the average loss of $\mathcal{Z}$ images according to Eq. (40) - Eq. (42), and update $\theta_{r}^{{\text{local}},n}$ according to SGD algorithm in Eq. (44);
@@
$\begin{split}\nabla_{\xi_{1}}\mathcal{J}_{\xi_{1}}=\nabla_{\xi_{1}}\Big{[}-\frac{1}{\mathcal{M}}\sum_{m=1}^{\mathcal{M}}{[r_{m}+\gamma\hat{Q}(s_{m+1},a_{m+1})}\\ \quad-Q_{\xi_{1}}(s_{m},a_{m})]^{2}\Big{]},\end{split}$ (53)
$\begin{split}\nabla_{\xi_{2}}\mathcal{J}_{\xi_{2}}=\nabla_{\xi_{2}}\Big{[}-\frac{1}{\mathcal{M}}\sum_{m=1}^{\mathcal{M}}{[r_{m}+\gamma\hat{Q}(s_{m+1},a_{m+1})}\\ \quad-Q_{\xi_{2}}(s_{m},a_{m})]^{2}\Big{]}.\end{split}$ (54)
@@ VI Simulation Results
We employ a modified ResNet-18, with a fixed dimension of 128-D, as the backbone model and utilize SGD with momentum as the optimizer. Momentum SGD is characterized by its strong randomness, rapid convergence, and ease of implementation, making it well-suited for optimizing the backbone model. Additionally, inspired by the concept of Cosine Annealing, we gradually reduce the learning rate at different stages of training to enhance the model’s training effectiveness.
Datasets: We utilize CIFAR-10 as the datasets, in which training datasets comprises 50,000 images distributed across 10 distinct categories, each of which consists 5,000 images. The testing datasets comprises 10000 images. We also establish two types of data distribution, including Independent and IID and Non-IID.
@@
(1) IID: Data obeying an IID are mutually independent and share the same distribution. Simultaneously, the statistical properties among data are similar, facilitating the generalization of the model. We uniformly allocate the 50,000 image data from CIFAR-10 to 95 training packages, ensuring that each training package has at least 520 images available. In Fig. 4, we show the distribution of image data categories on 10 training packages. As shown in Fig. 4(a), the quantity of each image data category on each vehicle is equal, indicating a uniform distribution state.
(2) Non-IID: IID offers theoretical convenience, but in practical scenarios, data satisfied with IID is rare. Therefore, it is beneficial to evaluate the model’s training performance under Non-IID conditions for transferring the model to real-world scenarios [66]. As depicted in Fig. 4(b) and 4(c), scenarios with Dirichlet distribution parameters $\alpha$ set to 0.1 and 1 are respectively shown. It can be observed that as $\alpha$ decreases, the gap between different categories increases, while a larger $\alpha$ makes the distribution closer to IID. We set the Dirichlet distribution parameter $\alpha$ to 0.1 to simulate Non-IID data in IoV, aiming to simulate situations where each training vehicle, due to limited perspectives and environmental constraints, collects images with uneven category distributions. Therefore, for CIFAR-10, we ensure that each training package has at least 515 Non-IID images.
Testing: We rank the predicted labels based on their probabilities from highest to lowest. If the most probable predicted label (i.e., the top label) matches the true label, the prediction is considered correct, and this is referred to as the Top1 accuracy. If the true label is among the top five predicted labels, the prediction is also considered correct, and this is referred to as the Top5 accuracy. For testing the performance of SAC algorithm, we employ the output $\varsigma^{\ast}$ to obtain $\pi^{\ast}(a_{k,t}|s_{k,t})$. Each experiment is conducted for three times, and the final result is the average of these three experiments.

### v2

@@ V DRL-BFSSL Algorithm
The optimization problem $\mathcal{C}$ is a NonLinear Programming (NLP) problem. As the number of the vehicles increases, it is hard to address rapidly with the traditional optimization algorithms. Thus, to achieve the optimal solution, we will consider a DRL offloading process. As we know, the SAC algorithm can deal with continuous variables in Non-convex function, and keep robustness and adaptability [56, 57]. Moreover, SAC is an off-policy DRL algorithm designed to enhance both the exploration of the policy and the system stability of the learning process. It achieves this by leveraging the maximum entropy framework, which encourages exploration by maximizing the entropy of the selected policy. This approach helps prevent premature convergence to suboptimal solutions. Additionally, SAC employs double Q-networks to mitigate the overestimation bias commonly encountered in value function approximation, further improving the model stability and reliability of the learning process [58, 59]. Thus, we adopt SAC algorithm to solve the optimization problem $\mathcal{C}$. In SAC algorithm, it consists of five Deep Neural Networks (DNNs), i.e., one actor network $\varsigma$, two critic networks $\xi_{1}$ and $\xi_{2}$, and two target critic networks $\varpi_{1}$ and $\varpi_{2}$. The two critic networks work to enhance network performance by minimizing positive bias. The two target networks have the same structure as the critic networks, aiming to improve training speed and stability. By utilizing reward function and Stochastic Gradient Descent (SGD) algorithm, SAC allows us to address the problem in continuous space.
@@ V-4 SAC-based solution
BS stores the parameters of a global model $\theta_{r-1}^{{\text{global}}}$ which is aggregated at the end of previous round $r-1$ and estimates the positions according to the uploaded velocities $\left\{v_{k,t}^{1},v_{k,t}^{2},\ldots v_{k,t}^{n}\ldots v_{k,t}^{N}\right\}$. Then BS can get state $s_{k,t}$ according to Eq. (34). Input $s_{k,t}$ into actor network and output $a_{k,t}$, including transmission powers and CPU frequencies for vehicles. Then, vehicle $n$ downloads the parameters of the global model $\theta_{r-1}^{{\text{global}}}$, transmission power $p_{k,t}^{n}$ , CPU frequency $f_{k,t}^{n}$ from the BS, and sets the parameter of global model to local model $\theta_{r}^{{\text{local},n}}$. If in the first round, the parameters of global model comes from $\theta_{0}^{{\text{global}}}$. If in the episode $k$, slot $1$, the parameters of SAC networks come from the output of episode $k-1$, slot $S^{\text{max}}$.
@@
Next, the maximum number of iterations $\mathcal{N}_{k,t}^{n}$ of local training for vehicle $n$ is calculated according to Eq. (16) and Eq. (17) based on the downloaded CPU frequency $f_{k,t}^{n}$. Then each training vehicle retrains the local model for $\mathcal{N}_{k,t}^{n}$ iterations. Specifically, for each iteration $\psi_{k,t}^{n}\in[1,\mathcal{N}_{k,t}^{n}]$ of local training, each vehicle $n$ encodes the images of the local data, where the encoding process is illustrated in Fig. 3. For each image $x_{r}^{i,n}\in\left\{x_{r}^{1,n},\ x_{r}^{2,n},\ \ldots x_{r}^{i,n},\ldots x_{r}^{\mathcal{Z},n}\right\}$ in vehicle $n$ undergoes two different data augmentation methods $\pi_{1}(\cdot)$ and $\pi_{2}(\cdot)$. $\pi_{1}(\cdot)$ performs a horizontal flip on the image with a 50% probability, followed by converting the image to grayscale with a 20% probability. $\pi_{2}(\cdot)$ randomly alters the image’s brightness, contrast, saturation, and hue with an 80% probability, followed by converting the image to grayscale with a 40% probability. It is also noted that $\pi_{1}(\cdot)$ and $\pi_{2}(\cdot)$ share the same original image when they process the image in different way [60]. We treat the remaining images, donated as $x_{r}^{j,n},j\in\left[1,\mathcal{Z}\right]\ \text{and}\ j\neq i$, as negative samples. Then, these two augmented images $\pi_{1}(x_{r}^{i,n})$ and $\pi_{2}(x_{r}^{i,n})$ along with $x_{r}^{j,n}$ are input into an $\chi$ function, which composed of an encoder with function $\mathcal{G}_{r}^{\text{local},n}$, and a project head including two dense neutral networks and a ReLU function. The $\chi$ function outputs anchor sample $q_{r}^{i,{n}}$, positive sample $k_{{r}}^{i,{n}}$, and encoded negative samples $k_{{r}}^{j,{n}}$, respectively, i.e.,
@@
Then, for each image, the aim is to minimize the loss function and get ideal optimize local mode $\hat{\theta}_{{r}}^{{\text{local}},n}$, i.e.,
@@
where $\theta_{{r}}^{{\text{local},n}}$ represents the parameters of local model of vehicle $n$ in round $r$. However, $\hat{\theta}_{{r}}^{{\text{local}},n}$ is an ideal value, and we use SGD to gradually bring $\theta_{{r}}^{{\text{local},n}}$ closer to $\hat{\theta}_{{r}}^{{\text{local}},n}$. The updating process can be described as follows:
@@
The issue of motion blur increases the difficulty of obtaining high-quality data while protecting privacy. To mitigate the negative impact of blurred images on model training, the proposed DRL-BFSSL algorithm enables to dynamically adjust the weight of blurred images in model aggregation. According to Eq. (46), we assign smaller weights to the parameters obtained from training vehicles with higher blur levels in the local model, thus reducing the impact of the motion blur and improving the performance of the global model. Then the BS employs a weighted federated algorithm to aggregate the parameters of $N$ models based on the blur level $\mathcal{B}_{r}^{n}$. The expression for the aggregated new global model is
@@
Repeat Step 2 to Step 4 until slot $t$ reaches to $S^{\text{max}}$. At this point, if the episode $k$ is divided by $K_{u}$, the five networks of SAC need to be updated, otherwise, it is not required. After that, episode $k$ moves to episode $k+1$, Slot $t$ reset to 00. Next, we will introduce the update process of five networks.
@@
The SAC algorithm prefers to achieve efficient and stable decision-making by continuously updating the policy through maximizing cumulative rewards and policy entropy. This is done by adding new data to the experience replay buffer and utilizing target networks. For each iteration of the updating process, randomly sample $\mathcal{M}$ tuples from the replay buffer to form a mini-batch for training. For the $m$-th tuple $(s_{m},a_{m},r_{m},s_{m+1})$ in the mini-batch, where $m\in[1,\mathcal{M}]$, inputting $s_{m}$ to the actor network to obtain $a_{m}^{\prime}$. It is worth noting that this action $a_{m}^{\prime}$ is not the same as $a_{m}$ in the tuple. Thus, the gradient of loss function of $\alpha$ can be calculated as
@@
1
2 Function UPDATE($\gamma$, $\alpha$, $\varsigma$, $\xi_{1}$, $\xi_{2}$, $\varpi_{1}$, $\varpi_{2}$, $\mathcal{M}$, $k$, $K_{t}$):
3
4      Randomly sample $\mathcal{M}$ tuples from replay buffer
5      Update $\alpha$ according to Eq. (50);
6      Update $\varsigma$ according to Eq. (52);
7      Update $\xi_{1}$ and $\xi_{2}$ according to Eq. (53) and Eq. (54);
8      if $k\mod K_{t}=0$ then
9             Update $\varpi_{1}$ and $\varpi_{2}$ according to Eq. (55) and Eq. (56);
10
@@ 4      Randomly sample $\mathcal{M}$ tuples from replay buffer
1 Initialize round $r\leftarrow 0$, episode $k\leftarrow 0$;
2 for episode $k$ from $1$ to $K^{\text{max}}$ do
3       slot $t\leftarrow 0$;
4       for slot $t$ from $1$ to $S^{\text{max}}$ do
5             $r\leftarrow(k-1)\cdot S^{\text{max}}+t$;
6             Obtain state $s_{k,t}$ based on distance and velocity;
7             Get action $a_{k,t}$ through SAC algorithm;
8             for $n=1,2,\dots,N$ in parallel do
9                   Download global model $\theta_{r-1}^{\text{global}}$, $p_{k,t}^{n}$, and $f_{k,t}^{n}$;
10                   $\theta_{r}^{{\text{local}},n}\leftarrow\text{LOCAL}(\theta_{r-1}^{\text{global}},p_{k,t}^{n},f_{k,t}^{n})$;
11
12            if $r=1$ then
13                   Vehicle $n\in[1,N]$ uploads the local model $\theta_{r}^{{\text{local}},n}$ and $v_{k,t}^{n}$ to BS with download transmission power $p_{k,t}^{n}$, along with velocity $v_{0}^{n}$;
14
15            else
16                   Vehicle $n\in[1,N]$ uploads the local model $\theta_{r}^{{\text{local}},n}$ and $v_{k,t}^{n}$ to BS with download transmission power $p_{k,t}^{n}$;
17
18            BS aggregates received local models based on $\mathcal{B}_{r}^{n}$ and updates $\theta_{r}^{\text{global}}$ according to Eq. (47);
19             Save $\theta_{r}^{\text{global}}$ to global model;
20
21      if episode $k$ mod $K_{u}$ = 0 then
22             UPDATE($\gamma$, $\alpha$, $\varsigma$, $\xi_{1}$, $\xi_{2}$, $\varpi_{1}$, $\varpi_{2}$);
23
24
25 Function Local($\theta_{r-1}^{\text{global}},p_{k,t}^{n},f_{k,t}^{n}$):
@@
26       Set the parameter of global model $\theta_{r-1}^{\text{global}}$ to local model $\theta_{r}^{{\text{local}},n}$;
27       Calculate iteration value $\mathcal{N}_{k,t}^{n}$ according to Eq. (16) and Eq. (17);
28       for each iteration $\psi_{k,t}^{n}$ from $1$ to $\mathcal{N}_{k,t}^{n}$ do
29             Prepare local model $\theta_{r}^{{\text{local}},n}$;
30             for image $i$ from $1$ to $\mathcal{Z}$ do
31                   Calculate $q_{r}^{i,n}$, $k_{r}^{i,n}$, and $k_{r}^{j,n}$ according to Eq. (38) - Eq. (40);
32
33            Calculate the average loss of $\mathcal{Z}$ images according to Eq. (41) - Eq. (43), and update $\theta_{r}^{{\text{local}},n}$ according to SGD algorithm in Eq. (45);
34
@@ 29             Prepare local model $\theta_{r}^{{\text{local}},n}$;
$\begin{split}\nabla_{\xi_{1}}\mathcal{J}_{\xi_{1}}=\nabla_{\xi_{1}}\Big{[}-\frac{1}{\mathcal{M}}\sum_{m=1}^{\mathcal{M}}{[r_{m}+\gamma\hat{Q}(s_{m+1},a_{m+1})}\\ \quad-Q_{\xi_{1}}(s_{m},a_{m})]^{2}\Big{]},\xi\in{\xi_{1},\xi_{2}}\end{split}$ (54)
@@ VI Simulation
$\varepsilon_{1}/\varepsilon_{2}/\varepsilon_{3}$ 0.3/0.3/0.4
We employ a modified ResNet-18, with a fixed dimension of $p_{j}$, as the backbone model and utilize SGD with momentum as the optimizer [65]. Momentum SGD is characterized by its strong randomness, rapid convergence, and ease of implementation, making it well-suited for optimizing the backbone model. Additionally, inspired by the concept of Cosine Annealing, we gradually reduce the learning rate at different stages of training to enhance the model’s training effectiveness.
VI-A Datasets
In intersection-related scenes, the categories of objects are relatively limited, and images of the same category are frequently collected. Therefore, CIFAR-10 is selected. In CIFAR-10, the training datasets comprises 50,000 images distributed across 10 distinct categories, each consisting of 5,000 images. The testing datasets comprises 10,000 images. We also establish two types of data distribution: IID, and Non-IID.
@@ VI-A Datasets
VI-A1 IID
Data obeying an IID are mutually independent and share the same distribution. Simultaneously, the statistical properties among data are similar, facilitating the generalization of the model. We uniformly allocate the 50,000 image data from CIFAR-10 to 95 training packages, ensuring that each training package has at least 520 images available. In Fig. 4, we show the distribution of image data categories on 10 training packages. As shown in Fig. 4(a), the quantity of each image data category on each vehicle is equal, indicating a uniform distribution state.
VI-A2 Non-IID
IID offers theoretical convenience, but in practical scenarios, data satisfied with IID is rare. Therefore, it is beneficial to evaluate the model’s training performance under Non-IID conditions for transferring the model to real-world scenarios [66]. As depicted in Fig. 4(b) and Fig. 4(c), scenarios with Dirichlet distribution parameters $\alpha$ set to 0.1 and 1 are respectively shown. It can be observed that as $\alpha$ decreases, the gap between different categories increases, while a larger $\alpha$ makes the distribution closer to IID. We set the Dirichlet distribution parameter $\alpha$ to 0.1 to simulate Non-IID data in IoV, aiming to simulate situations where each training vehicle, due to limited perspectives and environmental constraints, collects images with uneven category distributions.
VI-B Testing settings
We rank the predicted labels based on their probabilities from highest to lowest. If the most probable predicted label (i.e., the top label) matches the true label, the prediction is considered correct, and this is referred to as the Top1 accuracy. If the true label is among the top five predicted labels, the prediction is also considered correct, and this is referred to as the Top5 accuracy. For testing the performance of SAC algorithm, we employ the output $\varsigma^{\ast}$ to obtain $\pi^{\ast}(a_{k,t}|s_{k,t})$. Each simulation is conducted for three times, and the final result is the average of these three simulations.